Solve simultaneously
3x-2y=2
x²-3xy+xy²=129

Answer by Alan3354(69443)

3x-2y=2 ---> y = (3x-2)/2
x^2 - 3xy + xy^2 = 129
Sub for y
x^2 - 3x(3x-2)/2 + x((3x-2)/2)^2 = 129
x^2 - (9x^2-6x)/2 + x(9x^2-12x+4/4) = 129
4x^2 - 18x^2+ 12x + 9x^3 - 12x^2 + 4x = 516
9x^3 -26x^2 + 16x - 516 = 0
===========
No integer solutions.
1 real solution.
----------
Use iterative or graphical methods.